OpenStudy (anonymous):

If g(x) = 3(x − 2), find the value of x if g(x) = 6

OpenStudy (anonymous):

\[g(x) = 3(x-2)\] As g(x) = 6 \[6 = 3(x-2)\] From it try to find x..

OpenStudy (anonymous):

just tell me

OpenStudy (anonymous):

i dont no it

OpenStudy (anonymous):

we need to get the "3" to the left side of the equal sign

OpenStudy (anonymous):